Circular No. 6723

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 970616 P. J. Groot and T. J. Galama, University of Amsterdam; K. Hurley,

University of California at Berkeley; and C. Kouveliotou, Universities Space Research Association, on behalf of a larger collaboration, report: “Keck II images of GRB 970616, taken on June 21.6 and July 11.6 UT (unfiltered) and July 1.6 (R-band), show that the error box of the second ROSAT source, coinciding with the first ASCA variable source (IAUC 6687, 6722), contains four objects with (clockwise from north) magnitudes R = 24.8 ± 0.15 (object A), 25.9 ± 0.25 (B), 26.1 ± 0.25 (C), 25.8 ± 0.25 (D). For

the image, see http://www.astro.uva.nl/paulgr/grb/grb970616.html.

Located at the west and northeast edges, respectively, of the error box are two brighter objects, at R = 23.2 ± 0.1 (object E) and 21.7 ± 0.1 (F). None of these sources shows variation > 0.3 mag. Note also that the variability of the star reported by Galama et al. on IAUC 6687 reflects incorrect filter information for the June 21.6 observation; it shows no variability exceeding 0.1 mag between June 21 and July 11.”

GRB 970815 Groot et al. also write: “Observations of the error box of GRB 970815

(IAUC 6718, 6722) with the 1-m JKT reflector at La Palma on Aug. 16.10, 16.95, and 19.01 UT (V band) and the 3.5-m telescope at Apache Point Observatory on Aug. 16.3 and 16.5 (Sloan g′ band) show no sources with variability > 0.25 mag for V < 21.5.”

S. R. Kulkarni, S. C. Odewahn, and S. G. Djorgovski, California Institute of Technology; D. A. Frail, National Radio Astronomy Observatory; R. F. Goodrich, Keck Observatory; and L. M. Maxfield, University of California at Berkeley, report: “We observed the field of GRB 970815 in the Gunn i band using the CCD camera at the Palomar 1.52-m (P60) telescope during ∼ Aug. 18.19–18.35 UT. No statistically-significant variable sources changing by > 0.3 mag over the 4-hr interval were found, down to the limiting i-band magnitude of 20.5. Additional images were obtained at the Keck-II 10-m telescope during Aug. 20.26–20.29, reaching a limiting mag of I ∼ 24–24.5. No significantly-variable sources changing by > 0.5 mag over the period of 2 days were found, down to the limit of the P60 data. Very Large Array observations were made of the error box on Aug. 16.64, 17.10, and 21.13 at wavelengths of 6 and 3.6 cm. No reliable source detections were made inside this region above 0.35 mJy.”

1997 August 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green